OpenStudy (anonymous):

If anyone has done this in flvs or anything similiar, please help me! I will fan & give you a medal.

OpenStudy (anonymous):

Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. Using complete sentences, explain how to convert your birthday function into standard form. Create two additional quadratic functions, g(x) and h(x). The function g(x) will open the same direction as f(x), have the same vertex, but will be narrower. The function h(x) will open in the opposite direction as f(x), have the same vertex, but will be wider than f(x). Write your functions below and explain in complete sentences why those functions will meet the requirements in the question. Graph your functions. Include your graph below. Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. Explain the appropriate translations of your graphs, if you were born two months earlier and 5 days later. Use complete sentences.

OpenStudy (anonymous):

@slider1022 ..?

OpenStudy (anonymous):

think think THINK harder. You'll understand it eventually ^^

OpenStudy (anonymous):

My specialty is History not math sorry :(

OpenStudy (anonymous):

It's fine. Thanks though :)

OpenStudy (anonymous):

@AravindG ?

OpenStudy (anonymous):

you can cheat lol

OpenStudy (anonymous):

I just need help to figure it out.

OpenStudy (anonymous):

Alright lets see

OpenStudy (anonymous):

1) Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. General quadratic formula ... y = a(x-h)² + k choose a = 1 because it is easy to work with ... y = (x - 8)² + 15 2) Using complete sentences, explain how to convert your birthday function into standard form. Standard Form is ... Ax² + By = C So use FOIL on (x - 8)² and combine it with 15. Then put it in the Standard form above. 3) Create two additional quadratic functions, g(x) and h(x). A. The function g(x) will open the same direction as f(x), have the same vertex, but will be narrower. y = 4(x - 8)² + 15 B. The function h(x) will open in the opposite direction as f(x), have the same vertex, but will be wider than f(x). y = -0.1(x - 8)² + 15 Write your functions below and explain in complete sentences why those functions will meet the requirements in the question. The value of "a" is like a slope. The larger a, the narrower the slope. Also, if negate a, it will flip the graph. 4) Graph your functions. Include your graph below. Use your TI calculator or graph by hand. 5) Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. average rate of change = [f(7) - f(4)] / (7 - 4) It is simply the "slope of the line" between the two points. 6) Explain the appropriate translations of your graphs, if you were born two months earlier and 5 days later. Use complete sentences. y = (x - 6)² + 20 Will shift 2 units to the left and up 5 units.

OpenStudy (anonymous):

oooohthankksss:) my birthday is 02 09. how would i do it with my birthday?

OpenStudy (anonymous):

im not really sure haha sorry, and @shelbyygt520 why did you block me?

OpenStudy (anonymous):

Block you?

OpenStudy (anonymous):

yeah it said you blocked me

OpenStudy (anonymous):

I didn't block you lol

OpenStudy (anonymous):

I blocked sourwing lolll

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

thats so weird because when i tried to message you it says you blocked me

OpenStudy (anonymous):

beats me I have no idea I just checked it says under your name block not unblock

OpenStudy (anonymous):

GUYS LETS GET BACK ON TOPIC & TRY TO HELP KENDALL HERE OUT.

OpenStudy (anonymous):

whoops yeah sorry Kendall

OpenStudy (anonymous):

we need someone smarter in algebra 2 to help us lol

OpenStudy (anonymous):

@zpupster

OpenStudy (anonymous):

I've done FLVS what do you need? and yeah we do lol

OpenStudy (anonymous):

did you do alg 2?

OpenStudy (anonymous):

03.05 in seg. 1

OpenStudy (anonymous):

nope I'm one year to young sorry wish I could help

OpenStudy (anonymous):

aw girl lol it's okay. what math are you in?

OpenStudy (anonymous):

Geometry lol

OpenStudy (anonymous):

okay well if you ever need help, just ask me! :)

OpenStudy (anonymous):

will do! thanks :)

OpenStudy (anonymous):

@zpupster please help mee:((

OpenStudy (anonymous):

shelby message me

OpenStudy (anonymous):

whatt!! why did you block me? :((

OpenStudy (anonymous):

oh oh... looks like sourcing found ya @kendallnicole

OpenStudy (anonymous):

BECAUSE you weren't helping me.

OpenStudy (anonymous):

it wont let me message you because it says ive been blocked lmao

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

You don't want to block someone who is a math major LOL. That's ok, I still have Yahoo Answer :P

OpenStudy (anonymous):

THEN HELP ME PLEASEEEEE.

OpenStudy (anonymous):

I UNBLOCKED YOU.

OpenStudy (anonymous):

im just gonnna re-ask this question lol.

OpenStudy (anonymous):

yeah to much chitter chatter lol